# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v5

@@ Appendix A Summary of individual information
The problem in (19) serves purely as a theoretical intermediate step for establishing the regret upper bound in Theorem˜5.4. In practice, we do not compute $\hat{\psi}_{i,\bm{\theta}}$ for every $\bm{\theta}$. As discussed in Algorithm˜1, we only solve (19) once, with $\bm{\theta}=\tilde{\bm{\theta}}_{i}$, to obtain $\hat{\psi}_{i,\tilde{\bm{\theta}}i}$. Moreover, by point (4) in Proposition˜D.4, the function $\psi_{i,\bm{\theta}}$ is increasing (for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$). Since $h_{s_{i}}$ is also increasing, it follows that $\phi_{i,\bm{\theta}}$ inherits this monotonicity. Nonetheless, we did not include this monotonicity constraint in the definition of $\mathcal{S}_{i}$ in (19). This omission is justified because the theoretical convergence rate in Theorem˜D.6, and consequently the regret upper bound in Theorem˜5.4, is driven entirely by the concavity constraint, which imposes a higher-order smoothness condition than monotonicity. However, we do enforce monotonicity in the algorithmic implementation in Algorithm˜1.
@@
Let Assumptions 3.1, 3.4 and 5.1 hold. Then, by Proposition˜3.5 and Proposition˜D.4, $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave, where $s_{i}=c_{i}-1>-1$. Let $m=\pi n$, where $\pi=\int_{\mathcal{U}}f_{w}(u)du$ is the proportion of data points $\{w_{i}^{(t)}\}_{t\in[n]}$ lying in the set $\mathcal{U}$ as $n\rightarrow\infty$. For every $\gamma>4$ there exists $m_{0}$, a constant $\mathscr{C}_{i}>0$ and $\widetilde{\mathscr{C}}_{i}>0$ depending only on constants in the assumptions such that
@@
Proof of Theorem˜D.6.
We first prove the concentration inequality in equation 20.
@@
holds with probability at least $1-1/2(n+2)^{\gamma}$ for some $\gamma>0$, provided that $n$ is sufficiently large and $\epsilon_{n}\to 0$ as $n\to\infty$. Now, we apply Theorem˜H.5, using Remark˜H.6, with the following substitutions:
@@
where $h_{s_{i}}$ is defined in equation 4. Intersecting $\mathscr{Q}$ with the high-probability event in Theorem˜H.5 yields the desired result. The only remaining step is to verify that the assumptions of Theorem˜D.6 imply those of Theorem˜H.5.
@@
    $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave for some $s_{i}>-1$. According to Proposition˜D.4, this condition is satisfied provided $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$, which holds by Proposition˜3.5 and ˜3.4, with $s_{i}=c_{i}-1$.
@@
    ˜H.2 holds uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$ because the design points are drawn from an i.i.d. elliptically symmetric distribution with density generator $g>0$ and then $f_{\mathbf{p}}>0$ on any bounded interval. Thus, $\inf_{\bm{\theta}\in\mathbb{S}_{N-1}}\inf_{u\in\mathcal{U}}f_{\bm{\theta}^{\top}\mathbf{p}}(u)\geq C_{\mathscr{D}}$ for some $C_{\mathscr{D}}>0$.
@@
    ˜H.3 holds because $y_{i}^{(t)}$ are sub-gaussian (indeed $\psi_{i}$ is bounded and $\varepsilon_{i}^{(t)}$ are sub-gaussian) and $\psi_{i,\bm{\theta}}$ is bounded uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$ by Proposition˜D.4.
@@
    We prove that ˜H.4 holds uniformly in $\bm{\theta}$. Recall that $\phi_{i,\bm{\theta}}=d_{s_{i}}\circ\psi_{i,\bm{\theta}}$ (where $d_{s_{i}}\triangleq h^{-1}_{s_{i}}$ is defined in equation 4). Then $\psi_{i,\bm{\theta}}^{\prime}(u)=d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(u))\psi^{\prime}_{i,\bm{\theta}}(u)$. For every $u,v\in\mathcal{U}$ and $\bm{\theta}\in\mathbb{S}_{N-1}$ we have
@@
    The first component $|d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(u))-d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(v))|\leq L_{d^{\prime}}|u-v|$ by Section˜H.1, where $L_{d^{\prime}}=\sup_{x}d^{\prime\prime}_{s_{i}}(x)$ depends only on $s_{i}$ and bound of $\psi_{i}$. The component $|\psi^{\prime}_{i,\bm{\theta}}(u)|$ is bounded uniformly by Proposition˜D.4 point (4). The other component $|\psi^{\prime}_{i,\bm{\theta}}(u)-\psi^{\prime}_{i,\bm{\theta}}(v)|\leq L|u-v|$ for some $L>0$ independent of $\bm{\theta}$ by Proposition˜D.4 point (4) and $|d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(v))|$ is uniformly bounded by Section˜H.1. This proves that $\psi_{i,\bm{\theta}}^{\prime}$ is Lipschitz uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$.
To prove equation 22 we apply the triangular inequality
@@
and for the first component we apply the concentration inequality in equation 20 while for the second we use point (3) of Proposition˜D.4
@@
and then the concentration inequality in Proposition˜D.3, while retaining the dominating term $O((\log(m)/m)^{2/5})$. ∎
Appendix E Proof of Proposition˜D.3
@@
w.p. at least $1-2e^{-c\varsigma_{\min}^{2}n/16}$ (for $n$ sufficiently large) for certain constants $c>0$ that depends only on the variance proxy $\sigma_{x}$ of the sub-Gaussian random variable $\mathbf{x}^{(t)}$, and by Lemma˜E.2 the RHS satisfies
@@
with probability at least $1-\frac{2N}{n}$ (as long as $n\geq N$), where $\mathscr{B}_{i}^{**}$ is a constant defined in equation 23 that depends solely on $\lambda_{i}$ and the variance proxies of the sub-Gaussian random variables $y_{i}^{(t)}$ and $\mathbf{x}^{(t)}$. Putting the two inequalities together we get with probability at least $1-2e^{-c\varsigma_{\min}^{2}n/16}-\tfrac{2N}{n}$ that, for $n$ sufficiently large
@@
where $\mathscr{B}_{i}^{*}=2\frac{\mathscr{B}_{i}}{\varsigma_{\min}}$. Now we conclude. Recall that $\lambda_{i}>0$ where $\lambda_{i}$ is defined in equation 25. Let
@@
Proof of Lemma˜E.1.
To get a lower bound of the minimum eigenvalue of $\Sigma_{n}$ we use the following remark 5.40 of vershynin2010introduction: let $A$ be a $n\times N$ matrix whose rows $A_{i}$ are independent sub-gaussian random vectors in $\mathbb{R}^{N}$ with second moment matrix $\Sigma$ (non necessarily isotropic), then there exist $c,C>0$ depending only on the sub-gaussian norm $\sigma_{x}$, such that for every $r>0$ we have $\|\tfrac{1}{n}\sum_{i=1}^{n}A_{i}A_{i}^{\top}-\Sigma\|\leq\max\{\delta,\delta^{2}\}$ with probability at least $1-2e^{-cr^{2}}$, where $\delta=C\sqrt{\tfrac{N}{n}}+\tfrac{r}{\sqrt{n}}$. In our case, since $\Sigma_{n}$ is an average of $n$ i.i.d. random matrices with mean $\Sigma=\mathbb{E}\left[\mathbf{x}^{(t)}\mathbf{x}^{(t)\top}\right]$ and that $\left\{\mathbf{x}^{(t)}\right\}$ are sub-Gaussian random vectors (recall that $\mathbf{x}^{(t)}$ is sub-gaussian because the generator $g$ satisfies $g(x+y)=g(x)g(y)$ for all $x,y$), then there exist $c,C>0$ such that for each $r>0$ with probability at least $1-2e^{-cr^{2}}$,
@@
is a constant that depends solely on $\lambda_{i}$ (defined in equation 24) and the variance proxies $\sigma_{y_{i}}$ and $\sigma_{x}$ of the sub-Gaussian random variables $y_{i}^{(t)}$ and $\mathbf{x}^{(t)}$, respectively.
Proof of Lemma˜E.2.
@@
Using the known property of elliptically symmetric random variables (see, e.g. li1991sliced or balabdaoui2019least) we have that
@@
Appendix F Proof of Proposition˜D.4
@@ F.1 Proof of (1)
Then $\mathbf{y}^{\top}\mathbf{y}=y_{1}^{2}+\|\bm{\alpha}\|_{2}^{2}$. From the constraint in Equation˜27:
@@
where we defined the conditional center (the "shift")
@@ F.2 Proof of (2)
where we recall from Equation˜31 that $\mathbf{m}_{u}=\mathbf{m}+\Lambda\bm{\theta}\frac{\left(u-\bm{\theta}^{\top}\mathbf{m}\right)}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$. Note that
@@
where we used that $\bm{\theta_{i}},\bm{\theta}\in\mathbb{S}_{N-1}$. By equation F.1 we have
@@
proving that $\psi_{i,\bm{\theta}}\in\mathrm{C}^{2}(\mathcal{U})$ for all $\bm{\theta}$. By ˜3.1 we have that $0<\underline{B}_{\psi_{i}}\leq\psi_{i}\leq\bar{B}_{\psi_{i}}$, which immediately implies that $0<\underline{B}_{\psi_{i}}\leq\psi_{i,\bm{\theta}}\leq\bar{B}_{\psi_{i}}$ uniformly in $\bm{\theta}$, proving point (2).
@@ F.4 Proof of (4)
where $c(\bm{\theta})\triangleq\frac{d}{du}\bm{\theta}_{i}^{\top}\mathbf{m}_{u}=\frac{\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$ as defined in equation 33. Then we only need to prove that $\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}>0$ for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$. Recall by ˜5.1 that $\Lambda\in\mathbb{R}^{N\times N}$ is symmetric positive definite matrix with $c_{\min}\mathbb{I}\preccurlyeq\Lambda\preccurlyeq c_{\max}\mathbb{I}$ for some $0<c_{\min}\leq c_{\max}<\infty$, and that $\|\bm{\theta}\|_{2}=\|\bm{\theta}_{i}\|_{2}=1$. We are interested in the conditions under which
@@ F.5 Proof of (5)
is convex on $(0,\infty)\times(0,\infty)$, and then we can apply niculescu2006convex, which is an extension of the Jensen’s inequality for in $2$ variables. ∎
Appendix G Proof of Theorem˜5.4
@@
where we used ˜3.2: $0<\underline{B}_{\psi_{i}^{\prime}}\leq\psi_{i}^{\prime}\leq\bar{B}_{\psi_{i}^{\prime}}$ and $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ on $\mathcal{U}$ for some (unknown) $\underline{B}_{\psi_{i}^{\prime}},\bar{B}_{\psi_{i}^{\prime}},B_{\psi_{i}^{\prime\prime}}>0$. This implies
@@
If Assumptions in the Theorem hold, for $T$ sufficiently large we have $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]\lesssim\mathscr{X}_{i}\sqrt{N}(\tfrac{\log\tau}{\tau})^{2/5}$, where $\mathscr{X}_{i}=\max\left\{\mathscr{C}_{i},\mathscr{B}_{i}\right\}$, and $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are defined in equation 17 and equation 21, respectively. Moreover, there exists a unique value $\kappa_{i}^{\star}\in(0,1)$ that minimizes $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]$. This value satisfies the implicit equation:
@@
By Lemma˜G.1, Lemma˜G.2 have that for $t\geq\tau+1$ with $\tau$ sufficiently large,
@@
Note that we are retaining the dependence on $\mathscr{X}_{i}=\mathscr{X}_{i}(s_{i})$ to demonstrate the way that $s_{i}$ affects the rate of the regret. Substituting back the value $\tau\propto T^{\xi}$ into Appendix˜G, we get
@@
These bounds retain the constants $\mathscr{X}_{i}$ and $L_{\bm{\Gamma}}$ because they depend on the parameters $s_{1},\dots,s_{N}$ (or equivalently $c_{1},\dots,c_{N}$). Note that the dependence on $\{s_{j}\}_{j\in[N]}$ enters only through the multiplicative constant $\mathscr{X}_{i}$ and the additive constant $L_{\bm{\Gamma}}$, never through the powers of $T$. This makes explicit how the choice of $s_{i}$ influences the regret and the convergence to the NE. Below we show that if $s_{i}$ are the optimal-concave parameters (as defined in Definition˜L.3), i.e.
@@
    Multiplicative constants $\mathscr{X}_{i}$. The constant $\mathscr{X}_{i}=\{\mathscr{C}_{i},\mathscr{B}_{i}\}$ contains $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ (defined in equation 17–equation 21), which are the multiplicative constants in the concentration bounds for
@@
    respectively. Any value $s_{i}^{\prime}\neq s_{i}$ does not affect the rate of linear estimator $\widetilde{\bm{\theta}}_{i}$, but it could introduce a bias in the estimator $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$. Precisely, if we choose $s_{i}^{\prime}\leq s_{i}$, the inclusion property of Proposition˜L.1 implies that an $s_{i}$-optimal-concave function remains $s_{i}^{\prime}$-concave. Thus no bias is introduced in the concentration of $\|\psi_{i}-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}\|_{\infty}$, whereas a choice $s_{i}^{\prime}>s_{i}$ would. Then for $s_{i}^{\prime}\leq s_{i}$, since $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are uniformly bounded in $s_{i}$, i.e. $\mathscr{B}_{i},\mathscr{C}_{i}<C$ for some $C>0$ independent of $s_{i}$, they do not affect the regret or NE rates.
@@
G.1 Proof of Lemma˜G.1
Let $\tau$ be the minimum exploration phase among the $N$ firms and $t\in\{\tau+1,\tau+2,\dots,T\}$. Let $n_{i}^{(1)}$ be the length of the exploration phase to estimate $\bm{\theta}_{i}$ and $n_{i}^{(2)}$ the remaining length for estimating the link function $\psi_{i}$. Define the event $\mathcal{E}_{i}=\{\|\widetilde{\bm{\theta}}_{i}-\bm{\theta}_{i}\|\leq R_{n_{i}^{(1)}}\}$ where $R_{n_{i}^{(1)}}=\mathscr{B}_{i}\sqrt{\frac{N\log n_{i}^{(1)}}{n_{i}^{(1)}}}$, where $\mathscr{B}_{i}$ is defined in Proposition˜D.3, and
@@
By Proposition˜D.3 we have $\mathbb{E}[\mathscr{R}_{i}(p_{i}|\mathbf{p}_{-i})\mathbb{I}(\mathcal{E}_{i}^{c})]\lesssim\mathbb{P}(\mathcal{E}_{i}^{c})=Q_{n_{i}^{(1)}}=2e^{-c_{1}\varsigma_{\min}^{2}n_{i}^{(1)}/16}+\frac{2}{n_{i}^{(1)}}$.

### v6

@@ Appendix A Summary of individual information
The problem in (19) serves purely as a theoretical intermediate step for establishing the regret upper bound in Theorem 5.4. In practice, we do not compute $\hat{\psi}_{i,\bm{\theta}}$ for every $\bm{\theta}$. As discussed in Algorithm 1, we only solve (19) once, with $\bm{\theta}=\tilde{\bm{\theta}}_{i}$, to obtain $\hat{\psi}_{i,\tilde{\bm{\theta}}i}$. Moreover, by point (4) in Proposition D.4, the function $\psi_{i,\bm{\theta}}$ is increasing (for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$). Since $h_{s_{i}}$ is also increasing, it follows that $\phi_{i,\bm{\theta}}$ inherits this monotonicity. Nonetheless, we did not include this monotonicity constraint in the definition of $\mathcal{S}_{i}$ in (19). This omission is justified because the theoretical convergence rate in Theorem D.6, and consequently the regret upper bound in Theorem 5.4, is driven entirely by the concavity constraint, which imposes a higher-order smoothness condition than monotonicity. However, we do enforce monotonicity in the algorithmic implementation in Algorithm 1.
@@
Let Assumptions 3.1, 3.4 and 5.1 hold. Then, by Proposition 3.5 and Proposition D.4, $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave, where $s_{i}=c_{i}-1>-1$. Let $m=\pi n$, where $\pi=\int_{\mathcal{U}}f_{w}(u)du$ is the proportion of data points $\{w_{i}^{(t)}\}_{t\in[n]}$ lying in the set $\mathcal{U}$ as $n\rightarrow\infty$. For every $\gamma>4$ there exists $m_{0}$, a constant $\mathscr{C}_{i}>0$ and $\widetilde{\mathscr{C}}_{i}>0$ depending only on constants in the assumptions such that
@@
Proof of Theorem D.6.
We first prove the concentration inequality in (20).
@@
holds with probability at least $1-1/2(n+2)^{\gamma}$ for some $\gamma>0$, provided that $n$ is sufficiently large and $\epsilon_{n}\to 0$ as $n\to\infty$. Now, we apply Theorem H.5, using Remark H.6, with the following substitutions:
@@
where $h_{s_{i}}$ is defined in (4). Intersecting $\mathscr{Q}$ with the high-probability event in Theorem H.5 yields the desired result. The only remaining step is to verify that the assumptions of Theorem D.6 imply those of Theorem H.5.
@@
    $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave for some $s_{i}>-1$. According to Proposition D.4, this condition is satisfied provided $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$, which holds by Proposition 3.5 and 3.4, with $s_{i}=c_{i}-1$.
@@
    H.2 holds uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$ because the design points are drawn from an i.i.d. elliptically symmetric distribution with density generator $g>0$ and then $f_{\mathbf{p}}>0$ on any bounded interval. Thus, $\inf_{\bm{\theta}\in\mathbb{S}_{N-1}}\inf_{u\in\mathcal{U}}f_{\bm{\theta}^{\top}\mathbf{p}}(u)\geq C_{\mathscr{D}}$ for some $C_{\mathscr{D}}>0$.
@@
    H.3 holds because $y_{i}^{(t)}$ are sub-gaussian (indeed $\psi_{i}$ is bounded and $\varepsilon_{i}^{(t)}$ are sub-gaussian) and $\psi_{i,\bm{\theta}}$ is bounded uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$ by Proposition D.4.
@@
    We prove that H.4 holds uniformly in $\bm{\theta}$. Recall that $\phi_{i,\bm{\theta}}=d_{s_{i}}\circ\psi_{i,\bm{\theta}}$ (where $d_{s_{i}}\triangleq h^{-1}_{s_{i}}$ is defined in (4)). Then $\psi_{i,\bm{\theta}}^{\prime}(u)=d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(u))\psi^{\prime}_{i,\bm{\theta}}(u)$. For every $u,v\in\mathcal{U}$ and $\bm{\theta}\in\mathbb{S}_{N-1}$ we have
@@
    The first component $|d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(u))-d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(v))|\leq L_{d^{\prime}}|u-v|$ by Section H.1, where $L_{d^{\prime}}=\sup_{x}d^{\prime\prime}_{s_{i}}(x)$ depends only on $s_{i}$ and bound of $\psi_{i}$. The component $|\psi^{\prime}_{i,\bm{\theta}}(u)|$ is bounded uniformly by Proposition D.4 point (4). The other component $|\psi^{\prime}_{i,\bm{\theta}}(u)-\psi^{\prime}_{i,\bm{\theta}}(v)|\leq L|u-v|$ for some $L>0$ independent of $\bm{\theta}$ by Proposition D.4 point (4) and $|d_{s_{i}}^{\prime}(\psi_{i,\bm{\theta}}(v))|$ is uniformly bounded by Section H.1. This proves that $\psi_{i,\bm{\theta}}^{\prime}$ is Lipschitz uniformly in $\bm{\theta}\in\mathbb{S}_{N-1}$.
To prove (22) we apply the triangular inequality
@@
and for the first component we apply the concentration inequality in (20) while for the second we use point (3) of Proposition D.4
@@
and then the concentration inequality in Proposition D.3, while retaining the dominating term $O((\log(m)/m)^{2/5})$. ∎
Appendix E Proof of Proposition D.3
@@
w.p. at least $1-2e^{-c\varsigma_{\min}^{2}n/16}$ (for $n$ sufficiently large) for certain constants $c>0$ that depends only on the variance proxy $\sigma_{x}$ of the sub-Gaussian random variable $\mathbf{x}^{(t)}$, and by Lemma E.2 the RHS satisfies
@@
with probability at least $1-\frac{2N}{n}$ (as long as $n\geq N$), where $\mathscr{B}_{i}^{**}$ is a constant defined in (23) that depends solely on $\lambda_{i}$ and the variance proxies of the sub-Gaussian random variables $y_{i}^{(t)}$ and $\mathbf{x}^{(t)}$. Putting the two inequalities together we get with probability at least $1-2e^{-c\varsigma_{\min}^{2}n/16}-\tfrac{2N}{n}$ that, for $n$ sufficiently large
@@
where $\mathscr{B}_{i}^{*}=2\frac{\mathscr{B}_{i}}{\varsigma_{\min}}$. Now we conclude. Recall that $\lambda_{i}>0$ where $\lambda_{i}$ is defined in (25). Let
@@
Proof of Lemma E.1.
To get a lower bound of the minimum eigenvalue of $\Sigma_{n}$ we use the following remark 5.40 of Vershynin (2010): let $A$ be a $n\times N$ matrix whose rows $A_{i}$ are independent sub-gaussian random vectors in $\mathbb{R}^{N}$ with second moment matrix $\Sigma$ (non necessarily isotropic), then there exist $c,C>0$ depending only on the sub-gaussian norm $\sigma_{x}$, such that for every $r>0$ we have $\|\tfrac{1}{n}\sum_{i=1}^{n}A_{i}A_{i}^{\top}-\Sigma\|\leq\max\{\delta,\delta^{2}\}$ with probability at least $1-2e^{-cr^{2}}$, where $\delta=C\sqrt{\tfrac{N}{n}}+\tfrac{r}{\sqrt{n}}$. In our case, since $\Sigma_{n}$ is an average of $n$ i.i.d. random matrices with mean $\Sigma=\mathbb{E}\left[\mathbf{x}^{(t)}\mathbf{x}^{(t)\top}\right]$ and that $\left\{\mathbf{x}^{(t)}\right\}$ are sub-Gaussian random vectors (recall that $\mathbf{x}^{(t)}$ is sub-gaussian because the generator $g$ satisfies $g(x+y)=g(x)g(y)$ for all $x,y$), then there exist $c,C>0$ such that for each $r>0$ with probability at least $1-2e^{-cr^{2}}$,
@@
is a constant that depends solely on $\lambda_{i}$ (defined in (24)) and the variance proxies $\sigma_{y_{i}}$ and $\sigma_{x}$ of the sub-Gaussian random variables $y_{i}^{(t)}$ and $\mathbf{x}^{(t)}$, respectively.
Proof of Lemma E.2.
@@
Using the known property of elliptically symmetric random variables (see, e.g. Li (1991, page 319, comment following Condition 3.1) or Balabdaoui et al. (2019, Section 9.1)) we have that
@@
Appendix F Proof of Proposition D.4
@@ F.1 Proof of (1)
Then $\mathbf{y}^{\top}\mathbf{y}=y_{1}^{2}+\|\bm{\alpha}\|_{2}^{2}$. From the constraint in Equation 27:
@@
where we defined the conditional center (the ”shift”)
@@ F.2 Proof of (2)
where we recall from Equation 31 that $\mathbf{m}_{u}=\mathbf{m}+\Lambda\bm{\theta}\frac{\left(u-\bm{\theta}^{\top}\mathbf{m}\right)}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$. Note that
@@
where we used that $\bm{\theta_{i}},\bm{\theta}\in\mathbb{S}_{N-1}$. By (F.1) we have
@@
proving that $\psi_{i,\bm{\theta}}\in\mathrm{C}^{2}(\mathcal{U})$ for all $\bm{\theta}$. By 3.1 we have that $0<\underline{B}_{\psi_{i}}\leq\psi_{i}\leq\bar{B}_{\psi_{i}}$, which immediately implies that $0<\underline{B}_{\psi_{i}}\leq\psi_{i,\bm{\theta}}\leq\bar{B}_{\psi_{i}}$ uniformly in $\bm{\theta}$, proving point (2).
@@ F.4 Proof of (4)
where $c(\bm{\theta})\triangleq\frac{d}{du}\bm{\theta}_{i}^{\top}\mathbf{m}_{u}=\frac{\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$ as defined in (33). Then we only need to prove that $\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}>0$ for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$. Recall by 5.1 that $\Lambda\in\mathbb{R}^{N\times N}$ is symmetric positive definite matrix with $c_{\min}\mathbb{I}\preccurlyeq\Lambda\preccurlyeq c_{\max}\mathbb{I}$ for some $0<c_{\min}\leq c_{\max}<\infty$, and that $\|\bm{\theta}\|_{2}=\|\bm{\theta}_{i}\|_{2}=1$. We are interested in the conditions under which
@@ F.5 Proof of (5)
is convex on $(0,\infty)\times(0,\infty)$, and then we can apply Niculescu and Persson (2006, Theorem 3.5.3), which is an extension of the Jensen’s inequality for in $2$ variables. ∎
Appendix G Proof of Theorem 5.4
@@
where we used 3.2: $0<\underline{B}_{\psi_{i}^{\prime}}\leq\psi_{i}^{\prime}\leq\bar{B}_{\psi_{i}^{\prime}}$ and $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ on $\mathcal{U}$ for some (unknown) $\underline{B}_{\psi_{i}^{\prime}},\bar{B}_{\psi_{i}^{\prime}},B_{\psi_{i}^{\prime\prime}}>0$. This implies
@@
If Assumptions in the Theorem hold, for $T$ sufficiently large we have $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]\lesssim\mathscr{X}_{i}\sqrt{N}(\tfrac{\log\tau}{\tau})^{2/5}$, where $\mathscr{X}_{i}=\max\left\{\mathscr{C}_{i},\mathscr{B}_{i}\right\}$, and $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are defined in (17) and (21), respectively. Moreover, there exists a unique value $\kappa_{i}^{\star}\in(0,1)$ that minimizes $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]$. This value satisfies the implicit equation:
@@
By Lemma G.1, Lemma G.2 have that for $t\geq\tau+1$ with $\tau$ sufficiently large,
@@
Note that we are retaining the dependence on $\mathscr{X}_{i}=\mathscr{X}_{i}(s_{i})$ to demonstrate the way that $s_{i}$ affects the rate of the regret. Substituting back the value $\tau\propto T^{\xi}$ into Appendix G, we get
@@
These bounds retain the constants $\mathscr{X}_{i}$ and $L_{\bm{\Gamma}}$ because they depend on the parameters $s_{1},\dots,s_{N}$ (or equivalently $c_{1},\dots,c_{N}$). Note that the dependence on $\{s_{j}\}_{j\in[N]}$ enters only through the multiplicative constant $\mathscr{X}_{i}$ and the additive constant $L_{\bm{\Gamma}}$, never through the powers of $T$. This makes explicit how the choice of $s_{i}$ influences the regret and the convergence to the NE. Below we show that if $s_{i}$ are the optimal-concave parameters (as defined in Definition L.3), i.e.
@@
    Multiplicative constants $\mathscr{X}_{i}$. The constant $\mathscr{X}_{i}=\{\mathscr{C}_{i},\mathscr{B}_{i}\}$ contains $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ (defined in (17)–(21)), which are the multiplicative constants in the concentration bounds for
@@
    respectively. Any value $s_{i}^{\prime}\neq s_{i}$ does not affect the rate of linear estimator $\widetilde{\bm{\theta}}_{i}$, but it could introduce a bias in the estimator $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$. Precisely, if we choose $s_{i}^{\prime}\leq s_{i}$, the inclusion property of Proposition L.1 implies that an $s_{i}$-optimal-concave function remains $s_{i}^{\prime}$-concave. Thus no bias is introduced in the concentration of $\|\psi_{i}-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}\|_{\infty}$, whereas a choice $s_{i}^{\prime}>s_{i}$ would. Then for $s_{i}^{\prime}\leq s_{i}$, since $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are uniformly bounded in $s_{i}$, i.e. $\mathscr{B}_{i},\mathscr{C}_{i}<C$ for some $C>0$ independent of $s_{i}$, they do not affect the regret or NE rates.
@@
G.1 Proof of Lemma G.1
Let $\tau$ be the minimum exploration phase among the $N$ firms and $t\in\{\tau+1,\tau+2,\dots,T\}$. Let $n_{i}^{(1)}$ be the length of the exploration phase to estimate $\bm{\theta}_{i}$ and $n_{i}^{(2)}$ the remaining length for estimating the link function $\psi_{i}$. Define the event $\mathcal{E}_{i}=\{\|\widetilde{\bm{\theta}}_{i}-\bm{\theta}_{i}\|\leq R_{n_{i}^{(1)}}\}$ where $R_{n_{i}^{(1)}}=\mathscr{B}_{i}\sqrt{\frac{N\log n_{i}^{(1)}}{n_{i}^{(1)}}}$, where $\mathscr{B}_{i}$ is defined in Proposition D.3, and
@@
By Proposition D.3 we have $\mathbb{E}[\mathscr{R}_{i}(p_{i}|\mathbf{p}_{-i})\mathbb{I}(\mathcal{E}_{i}^{c})]\lesssim\mathbb{P}(\mathcal{E}_{i}^{c})=Q_{n_{i}^{(1)}}=2e^{-c_{1}\varsigma_{\min}^{2}n_{i}^{(1)}/16}+\frac{2}{n_{i}^{(1)}}$.